Exercise 9-90 (Algorithmic) (Appendix 9A) Calculating Bond Issue Price On January 1, 2020, University Theatres issued $461,00Exhibit A3.7 Future Value of a Single Amount FV=1(1+i) ni 1% 2% 3% 4% 5% 6% 7% 8% 9% 10% 12% 14% 16% 18% 20% 25% 30% 1 10100Exhibit A3.8 Present Value of a Single Amount PV= 7 (1+) Mi 1% 2% 3% 4% 5% 6% 7% 8% 9% 10% 12% 14% 16% 18% 20% 25% 30% + 0.99Exhibit A3.9 Future Value of an Al 2 1% 2% 3% 4% 5% 6% 7% 8% 9% 10% 12% 14% 16% 18% 20% 25% 30% 1 00000 100000 100000 100000Exhibit A3.10 (1+i) Present Value of an Annuity PVA = – m 1% 2% 3% 4% 5% 6% 7% 8% 9% 10% 12% 14% 18% 18% 20% 25% 30% 1 0.990
Answer #1

Bond face value = $461,000

Stated rate = 8%

Interest paid semi annually = $461,000 * 8% *6/12 = $18,440

tenure of bond (in years) = 15

tenure of bond (in periods) = 30

a) Market rate of interest = 6%

Market rate of interest = 6%/ 2 = 3% per period

Bond issue price = Face value* PV + Interest amount * PVA

Bond issue price = 461,000 * PV(3%,30) + 18,440 * PVA(3%,30)

Bond issue price = (461,000 * 0.41199) + (18,440 * 19.60044)

Bond issue price = $551,359.50

Bond issue price = $551,360

a) Market rate of interest = 10%

Market rate of interest = 10%/ 2 = 5% per period

Bond issue price = Face value* PV + Interest amount * PVA

Bond issue price = 461,000 * PV(5%,30) + 18,440 * PVA(5%,30)

Bond issue price = (461,000 * 0.23138) + (18,440 * 15.37245)

Bond issue price = $390,134.158

Bond issue price = $390,134